find the the area enclosed by y= x^2 , y= x+2 and x axis
 

Grade:12

1 Answers

Rajat
5 years ago
y= x + 2 and y = x^2 intersect at: (2,4) and (-1,1)
Area enclosed by y = X+2 and x axis from x=(-1) to X=2 is, 
\int_{-1}^{2} (x+2)dx = 15/2
Area enclosed by y= x^2 and x axis between X= (-1) and x=2
:\int_{-1}^{2} (x^2)dx = 8/3+1/3=3
Required area = (15/2)-3= 9/2